OpenStudy (anonymous):

solve for the variable. reduce the answer to lowest terms 1/8r=24

OpenStudy (shamim):

24*8r=1 right?

OpenStudy (anonymous):

I dunno, does it? I just need the answer

OpenStudy (anonymous):

1/8r is in fraction form

OpenStudy (shamim):

r=1/(8*24)=?